### v1

# Cosmological constraints from $N$-body simulations for future spectroscopic galaxy surveys at $2\leq z\leq 3$

Abstract
Determining the spatial curvature ($\Omega_{\rm k}$) independent of cosmic microwave background observations plays a key role in revealing the physics of the early universe. The Hubble tension is one of the most serious issues in modern cosmology. We investigate halo catalogs identified from $N$-body simulations at $z=2$ and 3, mimicking high-redshift galaxy surveys. We measure redshift-space correlation functions of halos from the two snapshots. We detect clear features of baryon acoustic oscillations and redshift-space distortions. We find that we can obtain a few percent constraints on both the geometric distances and growth of structure at the distant universe in future surveys. By taking into account the information of the underlying matter power spectrum, we demonstrate that we can also achieve constraint on the Hubble constant $H_{0}$ with a few percent as well as the spatial curvature with $|\Omega_{\rm k}|\lesssim 0.1$ by observing galaxies with the number density with $\bar{n}_{\rm g}\simeq 10^{-4}({\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}})$. Our analysis provides a timely forecast for the upcoming spectroscopic surveys, which target emission line galaxy or dusty star-forming galaxy samples.
I Introduction
Cosmological models have been tested via various observations including cosmic microwave background (CMB) Hinshaw et al. (2013); Planck Collaboration et al. (2016), gravitational lensing Abbott et al. (2018); Hikage et al. (2019), galaxy redshift surveys Ivanov et al. (2020); d’Amico et al. (2020); Kobayashi et al. (2022), Type-Ia supernovae Suzuki et al. (2012); DES Collaboration et al. (2024), etc. Cold dark matter with a cosmological constant, the $\Lambda$CDM model, has successfully explained all of these observations so far. However, there exist discrepancies between some parameters determined from the early universe and those from the local universe, known as the Hubble tension (Verde et al., 2019) and the $S_{8}$ tension Abdalla et al. (2022). The Dark Energy Spectroscopic Instrument (DESI) DESI Collaboration et al. (2024), the latest galaxy survey, has reported cosmological constraints that support dynamical dark energy with $w>-1$. However, combining them with other datasets still includes the cosmological constant ($w=-1$) within $2\sigma$ in a large cosmological parameter space (e.g., Choudhury and Okumura (2024)). Although the geometry of the universe has been found flat by the CMB, the flatness of the universe has not been constrained in the late universe observations (Takada and Doré, 2015; Terasawa et al., 2022, 2024). One way to tackle these issues is to extend the observations of cosmological objects to the more distant universe Okumura et al. (2016). In addition, cosmological spectroscopic experiments at $z\gtrsim 2$ have the potential to put meaningful constraints on curvature (Takada and Doré, 2015), neutrino mass (Takada et al., 2006; Dvorkin et al., 2019), inflation (Takada et al., 2006), and dark energy (Ferraro and Wilson, 2019).
Indeed, in the coming years a number of optical and near-infrared facilities in plan aim to conduct large-scale cosmological spectroscopic surveys for the $z\sim 2-4$ Universe, including the Subaru Prime Focus Spectrograph (PFS; Takada et al., 2014), DESI-II, Stage-V experiments such as Maunakea Spectroscopic Explorer (MSE; Marshall et al., 2019), MegaMapper (Schlegel et al., 2022), and SpecTel (Ellis and Dawson, 2019). In the longer wavelength regime, the 50-m single-dish submillimeter telescope AtLAST plans to perform the first cosmological spectroscopic surveys for over 100,000 galaxies in (sub)millimeter in degree square scales van Kampen et al. (2024).
The galaxy populations that these future facilities will target include Lyman Alpha Emitters (LAEs), Lyman Break Galaxies (LBGs), and dusty star-forming galaxies (DSFGs), and they are shown to probe halo masses that differ by up to two orders of magnitude, from $\sim$1011 M by the LAEs (Khostovan et al., 2019) to $\sim$1013 M by the DSFGs (Lim et al., 2020; Stach et al., 2021). The combined measurements of the various galaxy populations, therefore, allow us to suppress the cosmic variance via the multi-tracer technique McDonald and Seljak (2009); Ebina and White (2024).
In this paper, using dark matter $N$-body simulations, we aim to provide a first forecast on the cosmological constraints via clustering measurements that will be enabled by future spectroscopic surveys. As a first step, we focus on redshifts of 2 and 3, a natural extension of current state-of-the-art measurements DESI Collaboration et al. (2024). We measure correlation functions of dark-matter halos from the two snapshots. We then discuss how much cosmological information can be extracted from baryon acoustic oscillations (BAO) and redshift-space distortions (RSD) as well as the shape of the underlying matter power spectrum. Otherwise stated, we assume a spatially-flat $\Lambda$CDM model Planck Collaboration et al. (2016) as our fiducial cosmology throughout this paper.
II $N$-body Simulations and halo catalogs
We use $N$-body simulations run as part of the DARK QUEST project (Nishimichi et al., 2019). We employ $2048^{3}$ dark matter particles of mass $m_{p}=8.15875\times 10^{10}h^{-1}M_{\odot}$ in a cubic box on the side $2h^{-1}{\rm Gpc}$. We have the data set from three independent realizations and we specifically analyze clustering of dark matter halos of the snapshots at $z=2$ and $z=3$. Halos are identified using the ROCKSTAR algorithm (Behroozi et al., 2013). Their velocities and positions are determined by the average of the member particles within the innermost 10% of the subhalo radius. The halo mass, $M_{\rm h}$, is defined by a sphere with radius within which the enclosed average density is 200 times the mean matter density.
The statistical properties of halos that host galaxies at high redshifts, such as dusty galaxies and emission line galaxies, have not been accurately determined yet by observations. In this paper, thus, we do not consider any specific galaxy population but use all subhalos with some mass range. In order to see the halo mass dependence of the cosmological results, we adopt two mass ranges as shown in Table 1, referred to as the low- and high-mass samples, roughly corresponding to the mass ranges of emission line galaxies ($\sim 10^{-3}[{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}]$) and dusty galaxies ($\sim 10^{-4}[{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}]$), respectively.
Table 1: Properties of mock halo catalogs. The second column shows the mass range of the halos, $\bar{M}_{\rm h}$ is the average mass and $\bar{n}_{\rm g}$ is the number density.
$z$ $\log{M_{\rm h}/({\hbox{$~{}h^{-1}M_{\odot}$}})}$ $\log{\bar{M}_{\rm h}/({\hbox{$~{}h^{-1}M_{\odot}$}})}$ $10^{4}\bar{n}_{\rm g}$
(${}h^{3}{\rm~{}Mpc}^{-3}$)
2 $11.5-12.0$ 11.7 32.2
$12.5-13.5$ 12.8 4.05
3 $11.5-12.5$ 11.9 26.8
$12.5-13.5$ 12.7 1.02
III Redshift-space correlation functions
III.1 Measurements from $N$-body simulations
From the simulations, we measure the redshift-space correlation functions of halos, $\xi(r,\mu_{\bm{r}})$, where $r=|\bm{r}|$ with $\bm{r}$ being the separation vector between two points and $\mu_{\bm{r}}$ is the direction cosine between the line of sight and $\bm{r}$. We then determine the multipole moments of the correlation functions (Hamilton, 1992),
[…]
where ${\cal L}_{\ell}(\mu_{\bm{r}})$ is the $\ell$-th order Legendre polynomials.
We need to estimate a covariance matrix for the measured correlation functions for cosmological analysis. We adopt a bootstrap resampling method (e.g., Norberg et al., 2009),
$C_{ij}$ $\equiv C\left[\xi_{\ell}(r_{i}),\xi_{\ell^{\prime}}(r_{j})\right]$
$=\frac{1}{N_{\rm mock}-1}\sum^{N_{\rm mock}}_{k=1}\left[\xi_{\ell,k}(r_{i})-\bar{\xi}_{\ell}(r_{i})\right]$
×$\qquad\qquad\qquad\quad\times\left[\xi_{\ell,k}(r_{j})-\bar{\xi}_{\ell}(r_{j})\right],$ (2)
where $\bar{\xi}_{\ell}(r_{i})=N_{\rm mock}^{-1}\sum_{k=1}^{N_{\rm mock}}\xi_{\ell,k}(r_{i})$. We create fifty realizations from each simulation. Furthermore, in measuring the redshift-space density field, we regard each direction along the three axes of simulation boxes as the line of sight. Thus, we have $N_{\rm mock}=50\times 3\times 3=450$ realizations to construct the covariance matrix.
Figure 1: Multipoles of redshift-space correlation functions of halos at $z=2$ (left) and $z=3$ (right). The upper and lower rows show the results for the low- and high-mass halos, respectively. The red and blue points are the measured monopole and quadrupole moments from $N$-body simulations. The dashed curves are the model with the input cosmological parameters. The dotted curves are the best-fitting model at $60\leq r\leq 140{\hbox{$~{}h^{-1}$}{\rm~{}Mpc}}$ (see Sec. IV.2). Errors are the squared diagonal components of the covariance matrices.
Fig. 1 shows both the monopole and quadrupole moments of the redshift-distorted-space correlation functions of redshift 2 and 3 halos for the high-mass and low-mass bins. The error bars for the multipoles shown are the square root of the diagonal parts of the covariance matrix.
III.2 Linear theory
Here we present a theoretical model of the measured correlation functions to place cosmological constraints. Since theoretical models are derived in Fourier space, we first consider a model for the redshift-space power spectrum, $P(\mathbf{k})=P(k,\mu_{\bm{k}})$, with $k=|\bm{k}|$ and $\mu_{\bm{k}}$ the directional cosine. Given the underlying matter power spectrum, cosmological information is extracted from the dynamical and geometric distortions, which are probed, respectively, via redshift-space distortion (RSD) and Alcock-Paczynski (AP) effects (Alcock and Paczynski, 1979; Kaiser, 1987; Seo and Eisenstein, 2003; Matsubara, 2004; Okumura et al., 2008).
In this paper, we use the simplest model for RSD, the linear RSD model with a constant bias $b$, known as the Kaiser model (Kaiser, 1987). It is given by
$P(k,\mu_{\bm{k}})=(b+f\mu_{\mathbf{k}}^{2})^{2}P_{\mathrm{m}}(k),$ (3)
where $f(z)$ is the growth rate of the universe sensitive to modification of the gravity models and approximated by $f(z)=\Omega_{\rm m}^{6/11}(z)$ for the $\Lambda$CDM model. The BAO wiggles encoded in the underlying matter power spectrum in real space, $P_{\mathrm{m}}(k)$, is damped due to the non-linear structure formation. Since the effect is less significant at high redshift, we include it using a simple model (e.g, Chuang and Wang (2013)),
$P_{\mathrm{m}}(k)=P_{\mathrm{nw}}(k)+[P_{\mathrm{lin}}(k)-P_{\mathrm{nw}}(k)]\exp\left(-\frac{k^{2}}{2k_{*}^{2}}\right),$ (4)
where $P_{\mathrm{lin}}(k)$ and $P_{\mathrm{nw}}(k)$ are the linear matter power spectra with and without BAO wiggles, respectively Eisenstein and Hu (1998). The parameter $k_{*}$ controls the effect of the damping. Although it can be computed analytically Vlah et al. (2016), we simply treat it as a free parameter and marginalize it over to obtain cosmological constraints.
Given $P(k,\mu_{\bm{k}})$, the multipole moments are obtained as
[…]
Those with $\ell=0,2,4$ contain cosmological information and are given by $P_{\ell}(k)=K_{\ell}P_{m}(k)$, where Cole et al. (1994); Okumura and Jing (2011)
$(K_{0},K_{2},K_{4})=\left(b^{2}+\frac{2}{3}bf+\frac{1}{5}f^{2},\frac{4}{3}bf+\frac{4}{7}f^{2},\frac{8}{35}f^{2}\right).$ (6)
These multipole power spectrum can be converted into multipole correlation functions by
$\xi_{\ell}(r)=i^{\ell}\int\frac{dkk^{2}}{2\pi^{2}}P_{\ell}(k)j_{\ell}(kr),$ (7)
where $j_{\ell}$ is the $\ell$-th order spherical Bessel function.
To measure the correlation function, the halo position sampled with redshift and angular position need to be converted into the comoving positions by introducing a reference cosmology. The apparent mismatch between the reference and true cosmology causes a geometric distortion in the measured correlation function, known as the AP effect Alcock and Paczynski (1979). Taking this effect into account, the correlation functions presented above are modulated as (Xu et al., 2013)
$\xi_{0}(r)=\xi_{0}(\alpha r)+\frac{2}{5}\epsilon\left[3\xi_{2}(\alpha r)+\frac{d\xi_{2}(\alpha r)}{d\log(r)}\right],$ (8)
$\xi_{2}(r)=2\epsilon\frac{d\xi_{0}(\alpha r)}{d\log(r)}+\left(1+\frac{6}{7}\epsilon\right)\xi_{2}(\alpha r)+\frac{4}{7}\epsilon\frac{d\xi_{2}(\alpha r)}{d\log(r)}$
$\qquad\qquad\qquad\qquad\quad+\frac{4}{7}\epsilon\Bigg{[}5\xi_{4}(\alpha r)+\frac{d\xi_{4}(\alpha r)}{d\log(r)}\Bigg{]}.$ (9)
Here we introduced two parameters that characterize the isotropic and anisotropic shifts of the BAO position, $\alpha$ and $\epsilon$, respectively as (Eisenstein et al., 2005; Padmanabhan and White, 2008)
$\alpha=\Bigg{[}\frac{D_{\rm A}^{2}(z)}{D_{\rm A,fid}^{2}(z)}\frac{H_{\rm fid}(z)}{H(z)}\Bigg{]}^{1/3}\frac{r_{\rm s,fid}}{r_{\rm s}},$ (10)
$1+\epsilon=\Bigg{[}\frac{H_{\rm fid}(z)}{H(z)}\frac{D_{\rm A,fid}(z)}{D_{\rm A}(z)}\Bigg{]}^{1/3},$ (11)
where $H(z)$ and $D_{\rm A}(z)$ are the expansion rate and angular diameter distance, respectively, and $r_{\rm s}$ the sound horizon scale at recombination time. Quantities with the subscript $\rm{fid}$ are computed using fiducial cosmological parameters (see Sec. I). Detecting anisotropies of BAO enables us to separate the AP effect from the dynamical distortion effect.
In summary, given the underlying matter power spectrum, $P_{\mathrm{m}}(k)$, the measured redshift-space correlation function is characterized by a set of five parameters, $\theta=(b(z),f(z),H(z),D_{\rm A}(z),k_{*})$. The prediction with the input cosmological model for our simulations is shown by the black dashed curve in Fig. 1. Later in the paper, we perform direct constraints on cosmological parameters allowing the shape of the underlying power spectrum to vary, referred to as a full-shape analysis. For this case, the choice of the parameter space is arbitrary, and the above dynamical and geometric parameters are computed using the chosen parameters.
IV Cosmological analysis and results
In this section, we investigate how well one can constrain cosmological models at redshifts $z>2$ using our high-redshift simulations. In Sec. IV.2, we present constraints on the growth rate and expansion rate using dynamical and geometric distortions. In Sec. IV.3, we constrain cosmological parameters by taking into account the cosmological dependence of the underlying power spectrum shape.
IV.1 Setup for likelihood analysis
We perform the likelihood analysis using the measured correlation functions of halos with the theoretical models described in Sec. III.2. The $\chi^{2}$ statistic is given by
$\chi^{2}(\theta)=\sum^{2N_{\rm bin}}_{i=1}\sum^{2N_{\rm bin}}_{j=1}\Delta_{i}C_{ij}^{-1}\Delta_{j},$ (12)
where $\Delta_{i}\equiv\Delta(r_{i};\theta)=\xi_{\ell}^{\rm sim}(r_{i})-\xi_{\ell}^{\rm th}(r_{i};\theta)$ is the difference between the measured and predicted correlation functions with $\theta$ being a parameter set to be constrained. The analysis is performed on the adopted scales $r_{\rm min}\leq r_{i}\leq r_{\rm max}$ with the number of bins for each multipole being $N_{\rm bin}$.
Figure 2: Constraints on dynamical and geometric parameters, $(f(z),H(z),D_{\rm A}(z))$, obtained by the correlation functions of low-/high-mass halos at $60\leq r\leq 140{\hbox{$~{}h^{-1}$}{\rm~{}Mpc}}$. Nuisance parameters, $b$ and $k_{*}$, are marginalized over. The contours show the 68% and 95% C.L. from inward.
The correlation functions on small scales are affected by various non-linear effects, namely non-linear evolution, RSD, and biasing. On the other hand, on large scales the covariance matrix is biased due to our resampling method. Thus, we use a conservative range for the likelihood analysis and set $(r_{\rm rmin},r_{\rm max})=(60,140)[{\hbox{$~{}h^{-1}$}{\rm~{}Mpc}}\,]$. For this range, the non-linearity of the matter power spectrum is completely negligible. We therefore compute $P_{\mathrm{m}}(k)$ in linear theory using the public code CLASS (Blas et al., 2011). Since the nonlinear RSD also does not significantly affect our result, we can safely use the linear RSD model (Eq. 3). Although the nonlinear bias effect is severer as we consider relatively massive halos, we still adopt the linear bias for simplicity. We set the bin size to $5{\hbox{$~{}h^{-1}$}{\rm~{}Mpc}}$, and thus $N_{\rm bin}=16$. To perform a maximum likelihood analysis we use the Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013), under flat priors otherwise stated.
Figure 3: One-dimensional marginalized errors on $f(z)$ (upper panel) and $H(z)$ and $D_{\rm A}(z)$ (lower panel) as functions of redshift. Gray curves are the true values of parameters at redshift 2 and 3. Errors of 68% confidence level are shown, but those on $H$ and $D_{\rm A}$ are multiplied by 5 for illustration.
Cosmological constraints obtained below correspond to those expected from a galaxy survey of the volume of our simulations, $V=(2{\hbox{$~{}h^{-1}$}{\rm~{}Gpc}})^{3}$, smaller than typical upcoming surveys. However, since data in actual surveys are split into redshift bins for cosmological analysis, our constraints will roughly correspond to those from one redshift bin.
IV.2 Dynamical and geometric constraints
For the analysis of dynamical and geometric distortions, we have five parameters in total, $\theta=(b,f,H,D_{\rm A},k_{*})$, among which $b$ and $k_{*}$ are nuisance parameters that we want to marginalize over. The matter power spectrum is computed assuming the fiducial cosmology.
Table 2: Dynamical and geometric parameter constraints.
$z=2$ $z=3$
$\theta$ fiducial low mass high mass fiducial low mass high mass
$f$ 0.958 $0.903^{+0.017}_{-0.017}$ $0.868^{+0.085}_{-0.081}$ 0.981 $0.966^{+0.030}_{-0.030}$ $0.923^{+0.092}_{-0.091}$
$D_{\rm A}$ 1193.3 $1180.5^{+2.6}_{-2.6}$ $1189.1^{+5.8}_{-6.3}$ 1096.0 $1094.8^{+2.1}_{-2.1}$ $1085.9^{+6.3}_{-6.7}$
$H$ 204.09 $205.45^{+0.66}_{-0.67}$ $203.84^{+1.37}_{-1.33}$ 307.43 $306.46^{+0.92}_{-0.93}$ $308.75^{+2.65}_{-2.63}$
$b$ $1.932^{+0.016}_{-0.016}$ $3.394^{+0.104}_{-0.109}$ $3.226^{+0.027}_{-0.027}$ $5.651^{+0.113}_{-0.114}$
Fig. 2 shows two-dimensional error contours on the geometric and dynamical parameters for the four samples, normalized by their fiducial values. Fig. 3 shows the one-dimensional marginalized errors as a function of redshift. The constrained parameters are summarized in Table 2. The preliminary results of this analysis have been presented in (van Kampen et al., 2024). The best-fit model obtained here is shown by the dotted curves in Fig. 1.
Since the simulation box size is fixed, the strength of the constraints depends on the number density of halo samples. The best-fit models are consistent with the input (true) models of the simulations within $2\sigma$ level except for the low-mass sample at $z=2$. This offset is caused by the fact that our simple linear model is not very accurate for the precision measurement. However, such systematic effects can be easily controlled once a more sophisticated model is adopted to take account of nonlinear corrections (e.g., Nishimichi et al. (2020)). The precision of the growth rate is at most 10% for low number density and high mass halos. The precision is significantly improved to $\sim 2-3\%$ as the number density increases to a few times $10^{-3}{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}$. We see a similar trend for $H(z)$ and $D_{\rm A}(z)$, whose constraints improve from 0.8% to 0.3% and 0.5% to 0.2%, respectively. The Lyman-$\alpha$ survey result in the DESI Collaboration DESI Collaboration et al. (2024) gives $2\%$ and $2.4\%$ precision in $H(z)$ and $D_{\rm A}(z)$, respectively, with number density $n_{\rm g}=3\times 10^{-5}{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}$ at effective redshift 2.33. Our results at $z=2$ are, therefore, confirmed by observations.
IV.3 Full-shape constraints
Here we utilize the cosmological information encoded in the underlying matter power spectrum and directly constrain cosmological parameters. We consider a nonflat $\Lambda$CDM model with six parameters in total, $\theta=(b,~{}\Omega_{\rm m},~{}\Omega_{\rm k},~{}H_{0},~{}A_{\rm s},~{}k_{*})$, where $A_{\rm s}$ is the amplitude of the primordial curvature perturbations. We fix $\Omega_{\rm b}$ to the fiducial value because the baryon density is tightly constrained by CMB and big-bang nucleosynthesis Aver et al. (2013); Cooke et al. (2014); Planck Collaboration et al. (2016). We further fix the dark energy equation-of-state to a cosmological constant, $w_{0}=-1$, as the high-$z$ clustering is not sensitive to $w$. On the other hand, the high-redshift universe is a useful laboratory to test the flatness of the universe, $\Omega_{\rm k}$ (Takada and Doré, 2015).
Since there is a strong degeneracy between $b$ and $A_{\rm s}$ in linear theory, they cannot be tightly constrained Okumura et al. (2008). For sampling stability and efficiency of the MCMC, we adopt a Gaussian prior on $A_{\rm s}$ and marginalize it over together with the other nuisance parameters, $b$ and $k_{*}$. We confirmed that this does not affect constraints on the other parameters, since the degeneracies between $A_{\rm s}$ and other parameters are relatively small.111 Using the shape information of the power spectrum on nonlinear scales and the bispectrum enables one to simultaneously determine $b$ and the perturbation amplitude, and thus to investigate the $\sigma_{8}$ (or $S_{8}$) tension Philcox and Ivanov (2022). Such an investigation is left to our future work. Here we discuss only the constraints from high-mass halos, as our model of the correlation function is too simple to provide accurate cosmological constraints via the full-shape analysis from the low-mass halos.
Figure 4: Constraints on cosmological parameters $(\Omega_{\rm m},H_{0},\Omega_{\rm k})$ from the full-shape information of correlation functions of high-mass halo samples at $z=2$ and 3. Nuisance parameters, $b$ and $k_{*}$, as well as the amplitude $A_{\rm s}$ are marginalized over. The dotted lines indicate the fiducial values.
Fig. 4 shows the two-dimensional constraints on pairs of cosmological parameters. Table 3 summarizes the one-dimensional marginalized constraints. The best-fitting value of $b$ obtained here is fully consistent with that from the dynamical and geometric constraints. We obtain $1.3\%$ and $2.0\%$ constraints on $\Omega_{\rm m}$ from the clustering of halos at $z=2$ and $z=3$, respectively. Similarly, we achieve constraints of $1.7\%$ and $2.4\%$ in $H_{0}$. The reported discrepancy between the values of $H_{0}$ measured using early and late universe probes is about $5-10\%$. Our result thus demonstrates that achieving the number density of $\bar{n}_{\rm g}\simeq 10^{-4}({\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}})$ over the large volume at $z\geq 2$ provides a powerful test of the Hubble tension, independent of the local universe ($z\simeq 0$) and early universe ($z\simeq 1100)$ measurements.
Note that here our analyses show certain offsets for $\Omega_{\rm m}$ and $H_{0}$, while the results shown in the previous subsection are consistent with the input within $1\sigma$ level for these samples. These shifts are driven by the inaccuracy of our linear theory model with the simple BAO damping on small scales, where cosmological information can still be extracted. In fact, Fig. 1 shows that the best-fit model fails to capture behavior on smaller scales than the BAO scale, especially at $z=2$, while it agrees with the measurements up to the BAO scale. This issue can be easily restored once one employs a more sophisticated model and our main conclusion of constraining cosmology from high redshifts is unchanged.
Table 3: Cosmological parameter constraints.
$z=2$ $z=3$
$\theta$ fiducial high mass high mass
$\Omega_{\rm m}$ 0.3150 $0.3038^{+0.0042}_{-0.0042}$ $0.3017^{+0.0060}_{-0.0063}$
$H_{0}$ 67.3 $62.9^{+1.1}_{-1.0}$ $65.4^{+1.5}_{-1.5}$
$\Omega_{\rm k}$ 0 $0.023^{+0.092}_{-0.080}$ $0.004^{+0.135}_{-0.117}$
$10^{9}A_{\rm s}$ 2.12 $2.12^{+0.10}_{-0.10}$ $2.12^{+0.11}_{-0.11}$
$b$ $3.426^{+0.233}_{-0.221}$ $5.493^{+0.407}_{-0.376}$
We also obtain interesting constraints on the spatial curvature at both redshifts. Using the Fisher matrix formalism, Ref. Takada and Doré (2015) showed that one can achieve the precision of $|\Omega_{\rm k}|\lesssim 3\times 10^{-3}$ (68% C.L.) by observing galaxies with $10^{-4}({\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}})$ at $0\leq z\leq z_{\rm max}$ where $z_{\rm max}\simeq 2$, as the most ideal situation. With more realistic simulation-based analysis, we show that the galaxy clustering at one-single redshift of $z=2$ or $z=3$ provides a constraint on $\Omega_{\rm k}$, as $|\Omega_{\rm k}|\lesssim 0.1$, without relying on other observations such as the CMB.
V Summary and discussion
In this paper, we have studied the cosmological constraints expected from upcoming high-redshift galaxy surveys utilizing high-resolution $N$-body simulations at $z=2$ and 3. To obtain general conclusions, we did not consider specific galaxy populations or survey geometries but rather split our halo sample into low-mass and high-mass subsamples in a cubic box of side $2{\hbox{$~{}h^{-1}$}{\rm~{}Gpc}}$. We chose a conservative fitting range where a simple linear theory prediction recovered the input cosmological models. We considered two analysis methods, a template fitting for dynamical and geometric constraints and a full-shape analysis. With the former, we obtained $\sim 10\%$ constraint on $f(z)$ with the massive halo sample with $\bar{n}_{\rm g}\sim 10^{-4}{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}$ but the constraint improved to a few percent by increasing the number density by an order of magnitude. On the other hand, we achieved $<1\%$ constraints on $H(z)$ and $D_{\rm A}(z)$ even with the low-number density halo sample. Our full-shape analysis demonstrated that we can simultaneously investigate the Hubble tension and the flatness of the universe. Including the nonlinearity of the matter spectrum further enables us to address the $\sigma_{8}$ (or $S_{8}$) tension without combining the CMB priors.
Our analyses provide a timely forecast for the upcoming spectroscopic surveys. For example, the expected number density for the emission line galaxy (ELG) sample in PFS at redshift 2 is $2.7\times 10^{-4}{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}$. With already good precision in DESI’s results including galaxies spanning on a larger range of redshifts with a much lower number density of our samples, we can expect to see percent or sub-percent level constraints on cosmological parameters in PFS or future DESI releases that have higher number density. Our forecasts on massive halos are also helpful in guiding the design of the AtLAST spectroscopic surveys, which will predominantly target dusty star-forming galaxies at $z\gtrsim 2$. The expected number density obtained from AtLAST is comparable to those of both of our high-mass bins, which, based on our results, will allow for percent levels of constraints on cosmological parameters.
This paper did not consider any specific galaxy sample or survey geometry. For a more realistic cosmological forecast based on simulations, we need to generate light-cone output using multiple snapshots of $N$-body simulations (see e.g., Howlett et al. (2015); Ishikawa et al. (2024)). While the halo occupation distribution (HOD) of ELGs has been studied recently (e.g., Avila et al. (2020); Okumura et al. (2021)), that of dusty galaxies still contains several uncertainties. Thus, a cosmological forecast for dusty galaxies at high redshifts requires more effort. Such detailed investigations and analyses will be performed in our future work.
Acknowledgements.
We are grateful to the maintenance and administrative staff of our institutions, whose efforts in supporting our day-to-day work environment make our scientific discoveries possible. TO acknowledges support of the Taiwan National Science and Technology Council under Grants No. NSTC 112-2112-M-001-034- and NSTC 113-2112-M-001-011-. SYP and CCC acknowledge the support of the Taiwan National Science and Technology Council (111-2112M-001-045-MY3), as well as the Academia Sinica through the Career Development Award (AS-CDA-112-M02). TN acknowledges support by JSPS KAKENHI Grant Numbers JP20H05861, JP21H01081, JP22K03634, JP24H00215 and JP24H00221. KA is supported by Fostering Joint International Research (B) under Contract No. 21KK0050 and the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. JP24K17056. We acknowledge the use of ‘the HPC facility at ASIAA’ where part of the numerical analyses were done. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the Ministry of Science and Technology of Taiwan, and National Tsing Hua University.